OpenStudy (anonymous):

tracy can wallpaper 2 rooms in a new house in 4 hrs. together with her trainee they can wallpaper the 2 rooms in 3 hours. how long would it take the trainee working by herself to do the job?

OpenStudy (y2o2):

let tracy be x and the trainee be y: then \[2 \div (x+y) = 3\] \[2 \div x = 4\] therefore ; \[x+y = (2/3)\] \[x = 0.5\] \[y = (1/6)\] \[2 \div (y) = 2 \div (1/6) = 12\] The trainee would take 12 hours to finish the job alone.

OpenStudy (y2o2):

I think there's a simpler answer...but i don't know it :)